Calculate the velocity of an electron placed in the 2nd orbit of the Li2+ ion,. Also calculate the no. of revolutions per second that it makes around the nucleus. 

Grade:11

1 Answers

Swapnil Saxena
12 years ago

The velocity of an electorn in nth shell of the hydrogen like atom is ((2.18*10^6)*Z)/n. Putting Z=3(atomic no of lithium) and n=2,

=(2.18*10^6)*3/2

=(3.27*10^6)ms-1.

No of revolution per second 2(pi)r/v where r=(0.529*(n^2))/Z)*10-10

r=2(pi)((0.529*n^2)/Z)*10-10

r=2(pi)((0.529*4)/3)*10-10=(4.42*10-10)m

Time=(4.42*10-10)m/(3.27*10^6)ms-1=1.35*10-16   

No of revolutions is 1/T=.74*1016